The free energy of formation for C10H24(g) is 138.470 kJ/mol. What is the equilibrium constant at 25.0?C for its formation?
Answer #1

We know that the Gibbs free energy change \Delta G=-RTlnK

Where R =Universal gas constant, T is temperature in Deg.K and K is equilibrium constant

llnk=-\Delta G/RT

lnk= -138.47/(0.00831*298) =-138.47/2.4768 = -55.9163

K= 5.198X10-25